The shareholders' equity of Blue Corporation includes 500,000 of 1 par common stock and 100,000 of 6% cumulative preferred stock. The board of directors of Blue declared cash dividends of 4,000, 3000, and 24,000 as cash dividends in each of 2017, 2018 and 2019 respectively. What is the amount of dividends that common stockholders will receive in 2019?
Answer #1

Annual Preferred Dividend = 100000 x 6% = 6000
Preferred Dividend in arrears for 2017 = 6000 - 4000 = 2000
Preferred Dividend in arrears for 2018 = 6000 - 3000 = 3000

Dividend for Common Stockholders in 2019 = 24000 - 2000 - 3000 - 6000 = 13000

If you have any query, kindly comment with your query and please mark thumbs up.